OpenStudy (souvik):

A homogeneous ball is suspended from a thread with a length equal to the radius of the ball. How many times is the period of small oscillations of this pendulum greater than that of a mathematical pendulum suspended at the same distance from the centre of gravity?

OpenStudy (anonymous):

What is the meaning of mathematical pendulum - a point mass attached to a massless string?

OpenStudy (anonymous):

Hold on! Mathematically , time period of pendulum is \(\huge{T=2\pi \sqrt{\frac{I}{mgL}}}\) where I is the moment of inertia about the point of suspension and L is the distance b/w point of suspension and c.m. of the sphere. And if the size of sphere attached at the end is very small (mathematically zero), then \(I=mL^{2}\) This gives \(\huge{T=2\pi \sqrt{\frac{L}{g}}}\) However this time the sphere cannot be assumed as a point mass. So ,I will not be mL^2 Instead ,we have to use parallel axis theorem and calculate the moment of inertia about the point of suspension. On the left is our given pendulum. \(\huge{I=m(2R)^{2}+\frac{2}{5}mR^{2}=\frac{22}{5}mR^{2}}\) This is the value of I for the given pendulum. So its time period will be(as given by first formula) will be, \( \huge{T=2\pi \sqrt {\frac{11R}{5g}}}\) And the time period for mathematical pendulum will be as given by second formula \(\huge{T=2\pi \sqrt{\frac{2R}{g}}}\)

OpenStudy (souvik):

so the time period of this pendulum/that of mathematical pendulum= \(\sqrt{11/10}\)~1.05 yeah i got it...thanks..:)